# Twenty-four thousand hours of GREENBURST observations with the GBT

### v1

@@ 3 Results
Figure 5: Example detections of individual pulses from the three previously known FRBs (FRB 20121102A, FRB 20190520B, and FRB 20200120E) observed during follow-up observations of these sources as well as the discovery pulse for GBP 220718. In each case the top panel shows the dedispersed pulse, the middle panel shows the ‘waterfall’ plot of frequency vs time, while the bottom panel shows the trial DM as a function of time.
@@ 4.1.1 Nulling analysis
We checked this value of NF using the earlier method of Ritchings (1976), who generated histograms of fluence from the on-pulse region and a subset of the off-pulse region, and computed the scaling factor required to make the two comparable; they identify this scaling factor with the nulling fraction. This results in a ${\rm NF}=80\pm 7\%$, consistent with the value found above.
Table 2: Observed and derived parameters for PSR J0039+5407
@@
DM-derived distance, $d$ (kpc) 2.4
Height above the Galactic plane, $z$ (pc) 360
@@ 4.2.2 Dispersion measure of the main pulse
As shown in Fig. 5, the narrow-band nature of GBP 220718 leads to a large bow-tie feature in the DM-time plane that is consistent with a pulse with a DM of 145.5 cm-3 pc, substantially higher than the Galactic contribution predicted by NE2001 (Cordes and Lazio, 2002) or YWM16 (Yao et al., 2017) which predict DMs of 65.7 and 71.9 cm-3 pc, respectively. It is possible that GBP 220718 is actually within the Milky Way, and that the apparent DM excess is due to inaccuracies in Galactic DM models along the line of sight. We searched the ATNF catalog for any pulsars with DMs exceeding the NE2001 maximum line-of-sight DM by 79.8 cm-3 pc or the YMW16 maximum line-of-sight DM by 73.6 cm-3 pc, the excesses seen from this FRB. For each model, only $\sim$0.6% of pulsars showed such large excesses, making this an unlikely explanation. In fact, most of that group of pulsars lie in globular clusters or the Magellanic Clouds, which explain many of the discrepancies. Additionally, pulsars near GBP 220718 in the sky do not appear to have anomalously high dispersion measures, and the DM of GBP 220718 is at least four times that of every known pulsar within 10 degrees of its sky position.
If GBP 220718 is astrophysical, a simple estimate of the implied redshift, $z$, comes from the fact that extragalactic DM scales roughly as $1000z$ (see, e.g., Petroff et al., 2019). This results in an estimated redshift for GBP 220718 of $z\sim 0.07$. Care should be taken here, however, since this calculation does not account for the free electrons in the Galactic halo which should also add to the source’s DM, and is capable of contributing well over 100 cm-3 pc along some lines of sight (Price et al., 2021). In the direction of GBP 220718, the halo model of Yamasaki and Totani (2020) predicts a contribution of only 32 cm-3 pc to the DM of GBP 220718, which is not enough to bridge the gap between the observed DM and the maximum DM in its direction predicted by electron density models.
@@ 5 Conclusions
The first 20,000 hours of GREENBURST observations resulted in the detections of 50 pulsars, one of which (PSR J0039+5407) was previously unknown, three previously known FRBs and one dispersed pulse of unknown origin (GBP 220718). PSR J0039+5407 is a 2.2 s pulsar with a nulling fraction on the order of 80%. GBP 220718 was originally detected as a narrow-band signal with a DM of 145.5 cm-3 pc. Based on FRB observations to date, such properties are consistent with a nearby repeating FRB. However, in spite of significant follow-up observations with other telescopes, no additional events were found. In addition, upon closer inspection of the data taken around the epoch of detection for GBP 220718, however, we found a number of additional pulses that are consistent with zero DM signals in the same part of the band. We are faced with a dilemma, however, in that the higher order moments of GBP 220718 are inconsistent with most known sources of terrestrial origin. Further observations are necessary to make a firm conclusion as to whether this source is astrophysical or terrestrial in nature.
@@ Acknowledgements
The Green Bank Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. We thank Green Bank Observatory staff Ryan Lynch and Evan Smith for their assistance with the followup observations. GREENBURST operations have been supported by NSF Astrononomy and Astrophysics awards 1616042 and 2406570. S.P. and D.R.L. acknowledge support from NSF award 2406570. S.T. and D.R.L. acknowledge support from NSF award 2307581. S. S. is a WVU Ruby Doctoral Fellow. M. F. acknowledges support from the WV Spacegrant Consortium. F.A.D is supported by an NRAO Jansky Fellowship. We acknowledge that CHIME is located on the traditional, ancestral, and unceded territory of the Syilx/Okanagan people. We are grateful to the staff of the Dominion Radio Astrophysical Observatory, which is operated by the National Research Council of Canada. CHIME operations are funded by a grant from the NSERC Alliance Program and by support from McGill University, University of British Columbia, and University of Toronto. CHIME was funded by a grant from the Canada Foundation for Innovation (CFI) 2012 Leading Edge Fund (Project 31170) and by contributions from the provinces of British Columbia, Québec and Ontario. The CHIME/FRB Project, which enabled development in common with the CHIME/Pulsar instrument, was funded by a grant from the CFI 2015 Innovation Fund (Project 33213) and by contributions from the provinces of British Columbia and Québec, and by the Dunlap Institute for Astronomy and Astrophysics at the University of Toronto. Additional support was provided by the Canadian Institute for Advanced Research (CIFAR), the Trottier Space Institute at McGill University, and the University of British Columbia. The CHIME/Pulsar instrument hardware is funded by the Natural Sciences and Engineering Research Council (NSERC) Research Tools and Instruments (RTI-1) grant EQPEQ 458893-2014.

### v2

@@ 3 Results
Figure 5: Example detections of individual pulses from the three previously known FRBs (FRB 20121102A, FRB 20190520B, and FRB 20200120E) observed during follow-up observations of these sources as well as the discovery pulse for GBP 220718. In each case the top panel shows the dedispersed pulse, the middle panel shows the ‘waterfall’ plot of frequency vs time, while the bottom panel shows the trial DM as a function of time. In each of the data visualizations, the median has been subtracted from the data and the resulting arrays have been divided by their respective standard deviation.
@@ 4.1.1 Nulling analysis
We checked this value of NF using the earlier method of Ritchings (1976), who generated histograms of fluence from the on-pulse region and a subset of the off-pulse region, and computed the scaling factor required to make the two comparable; they identify this scaling factor with the nulling fraction. This results in a ${\rm NF}=80\pm 7\%$, consistent with the value found above. Further observations of this pulsar to obtain high-quality pulse sequences to better characterize the single-pulse behaviour are strongly encouraged.
Table 2: Observed and derived parameters for PSR J0039+5407. The distance parameters were derived using the NE2025 model (Ocker and Cordes, 2026).
@@
DM-derived distance, $d$ (kpc) 3.0
Height above the Galactic plane, $z$ (pc) 450
@@ 4.2 GBP 220718
4.2.1 Search for low earth orbiting spacecraft
As the number of low-Earth orbit (LEO) satellites continues to increase (McDowell, 2020), distinguishing between transient RFI and genuine astrophysical signals has become vital. Indeed, a very narrow (ns duration) bright radio pulse from a decomissioned LEO satellite was recently observed in the 700–1000 MHz band with ASKAP (James et al., 2025). We conducted a search for LEO satellites above the Green Bank horizon at the time of GBP 220718. Using historical two-line element data from the Space-Track repository (United States Space Command, 2022), we reconstructed the distribution of LEOs and found none within $5^{\circ}$ of the GBT boresight (Azimuth: $45.3^{\circ}$, Elevation: $77.5^{\circ}$). The closest LEO at the time of the event was COSMOS 1320, launched in 1981, at an angular separation of $7.2^{\circ}$, while the nearest Starlink satellite (STARLINK-3843, a version 1.5 spacecraft) was over $17^{\circ}$ away. While Starlink V1.5 satellites are designed for Ku-band communication, they emit unintended electromagnetic radiation in the 110–188 MHz range (Di Vruno et al., 2023; Grigg et al., 2025). Whether this leakage extends into L-band remains uncharacterized. However, given that these satellites would be in the far field of the GBT at these frequencies (where the beam is well-defined and sidelobe response is significantly attenuated), and given the large angular separations observed here, we consider it unlikely that LEO satellites produced GBP 220718.
@@ 4.2.3 Dispersion measure of the main pulse
As shown in Fig. 5, the narrow-band nature of GBP 220718 leads to a large bow-tie feature in the DM-time plane that is consistent with a pulse with a DM of 145.5 cm-3 pc, substantially higher than the Galactic contribution predicted by NE2025 (Ocker and Cordes, 2026) or YWM16 (Yao et al., 2017) which predict DMs of 55.2 and 71.9 cm-3 pc, respectively. It is possible that GBP 220718 is actually within the Milky Way, and that the apparent DM excess is due to inaccuracies in Galactic DM models along the line of sight. We searched the ATNF catalog for any pulsars with DMs exceeding the NE2025 maximum line-of-sight DM by 90.3 cm-3 pc or the YMW16 maximum line-of-sight DM by 73.6 cm-3 pc, the excesses seen from this FRB. For each model, <1<1% of pulsars showed such large excesses, making this an unlikely explanation. In fact, most of that group of pulsars lie in globular clusters or the Magellanic Clouds, which explain many of the discrepancies. Additionally, pulsars near GBP 220718 in the sky do not appear to have anomalously high dispersion measures, and the DM of GBP 220718 is at least four times that of every known pulsar within 10 degrees of its sky position.
If GBP 220718 is astrophysical, we note that it is spatially coincident with the galaxy LEDA 2277498 (Paturel et al., 2003). This candidate host, also identified as SDSS J072513.14+463143.5, lies at an angular separation of approximately $17^{\prime\prime}$ from our best-fit position, well within the estimated $1\sigma$ uncertainty. The galaxy is a faint dwarf system with a photometric redshift of $z_{\text{phot}}\approx 0.082$ (Ahumada et al., 2020). The extragalactic dispersion measure ($\text{DM}_{\text{ext}}$) of the burst further supports this association. While a crude $\text{DM}_{\text{ext}}\sim 1000z$ scaling (Petroff et al., 2019) suggests $z\sim 0.07$, a more rigorous accounting of the foreground is required. In the direction of GBP 220718 ($l=176.3^{\circ},b=23.4^{\circ}$), the NE2025 model predicts a Galactic disk contribution of $\text{DM}_{\text{ISM}}=38$ pc cm-3. Accounting for a halo contribution of $\sim 32$ pc cm-3 (Yamasaki and Totani, 2020) and an expected IGM contribution of $\text{DM}_{\text{IGM}}\approx 66$ pc cm-3 at $z=0.082$ (Macquart et al., 2020), the observed DM is broadly consistent with an origin in LEDA 2277498. While these model-dependent subtractions suggest a low host contribution ($\text{DM}_{\text{host}}\lesssim 30$ pc cm-3), the large uncertainties in the halo and IGM components preclude a precise measurement. Nevertheless, the alignment in DM-space supports the possibility that the progenitor is located in a low-density environment, perhaps in the outskirts of the galaxy.
@@ 5 Conclusions
The first 20,000 hours of GREENBURST observations resulted in the detections of 50 pulsars, one of which (PSR J0039+5407) was previously unknown, three previously known FRBs and one dispersed pulse of unknown origin (GBP 220718). PSR J0039+5407 is a 2.2 s pulsar with a nulling fraction on the order of 80%. GBP 220718 was originally detected as a narrow-band signal with a DM of 145.5 cm-3 pc. Based on FRB observations to date, such properties are consistent with a nearby repeating FRB. However, in spite of significant follow-up observations with other telescopes, no additional events were found. In addition, upon closer inspection of the data taken around the epoch of detection for GBP 220718, however, we found a number of additional pulses that are consistent with zero DM signals in the same part of the band. We searched for possible LEO satellites above the horizon at the time of the observation of GBP 220718, but found none. In addition, the higher order moments of GBP 220718 are inconsistent with most known sources of terrestrial origin. If astrophysical, GBP 220718 appears to be coincident with the nearby galaxy LEDA 2277498. Further observations are necessary to make a firm conclusion as to whether GBP 220718 is astrophysical or terrestrial in nature.
@@ Acknowledgements
The Green Bank Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. We thank Green Bank Observatory staff Ryan Lynch and Evan Smith for their assistance with the followup observations. GREENBURST operations have been supported by NSF Astrononomy and Astrophysics awards 1616042 and 2406570. S.P. and D.R.L. acknowledge support from NSF award 2406570. S.T. and D.R.L. acknowledge support from NSF award 2307581. S. S. and K. E. H. are WVU Ruby Doctoral Fellows. M. F. acknowledges support from the WV Spacegrant Consortium. E. J. M. acknowledges NSF support from the WVU Research Experiences for Undergraduates program under award 2348764. F.A.D is supported by an NRAO Jansky Fellowship. We acknowledge that CHIME is located on the traditional, ancestral, and unceded territory of the Syilx/Okanagan people. We are grateful to the staff of the Dominion Radio Astrophysical Observatory, which is operated by the National Research Council of Canada. CHIME operations are funded by a grant from the NSERC Alliance Program and by support from McGill University, University of British Columbia, and University of Toronto. CHIME was funded by a grant from the Canada Foundation for Innovation (CFI) 2012 Leading Edge Fund (Project 31170) and by contributions from the provinces of British Columbia, Québec and Ontario. The CHIME/FRB Project, which enabled development in common with the CHIME/Pulsar instrument, was funded by a grant from the CFI 2015 Innovation Fund (Project 33213) and by contributions from the provinces of British Columbia and Québec, and by the Dunlap Institute for Astronomy and Astrophysics at the University of Toronto. Additional support was provided by the Canadian Institute for Advanced Research (CIFAR), the Trottier Space Institute at McGill University, and the University of British Columbia. The CHIME/Pulsar instrument hardware is funded by the Natural Sciences and Engineering Research Council (NSERC) Research Tools and Instruments (RTI-1) grant EQPEQ 458893-2014. We thank Matthew Stevenson for assistance with the LEO satellite analysis and the referee for very useful comments on an earlier version of the manuscript.